OpenStudy (ktbug3):

will fan n medal What is the circumference of a 12-inch diameter pizza? Leave the answer in terms of π. A) 3π B) 6π C) 12π D) 24π

OpenStudy (ktbug3):

c=pie*d

OpenStudy (ktbug3):

hey babby since u like math can u help @cmtboy2

OpenStudy (cmtboy2):

12tt

OpenStudy (ktbug3):

ok thnx love ya i got more

OpenStudy (ktbug3):

Calculate the area of a circle with a radius of 2 cm and a circle with a radius of 4 cm. Leave your answers in terms of pi. When the radius doubles, does the area of the circle double? A) 4π cm2; 8π cm2; NO B) 4π cm2; 16π cm2; NO C) 4π cm2; 16π cm2; YES D) 4π cm2; 8π cm2; YES

OpenStudy (ktbug3):

c=tt*d

OpenStudy (faiqraees):

c

OpenStudy (cmtboy2):

c

OpenStudy (ktbug3):

thnx

OpenStudy (ktbug3):

no it was b

OpenStudy (cmtboy2):

it was?

OpenStudy (ktbug3):

yeah

OpenStudy (cmtboy2):

oh sorry

OpenStudy (ktbug3):

its fine

OpenStudy (cmtboy2):

your turn help me lol

OpenStudy (ktbug3):

ok

OpenStudy (ktbug3):

tag me

OpenStudy (ktbug3):

π is the ratio of _______________ to _______________. A) area to diameter B) perimeter to area C) radius to diameter D) circumference to diameter Submit

OpenStudy (anonymous):

circumference= pi*d= 12*pi

OpenStudy (cmtboy2):

area to diameter

OpenStudy (ktbug3):

k thnx

OpenStudy (cmtboy2):

what did you get

OpenStudy (ktbug3):

wait...................it was d....

OpenStudy (cmtboy2):

no its a

OpenStudy (ktbug3):

its said it was wrong that it was d....

OpenStudy (ktbug3):

lol